# IRIF, Université Paris Cité, France

### v1

@@ 5.1 Evaluation Metrics
We begin by briefly introducing the metrics used to evaluate the effectiveness of our improved signature extraction:
  • Weight Recovery Rate denotes the proportion of weights successfully recovered in a layer, including those that can be safely set to zero without impacting the extraction, relative to the total number of weights in that layer. We refer to the other weights as missed weights.
  • Layer Coverage refers to the percentage of the input space where no unrecovered weight or component of the layer is active.
  • Model Coverage refers to the percentage of the input space where no unrecovered weight or component of the whole model is active.
Not all weights contribute equally to the network’s behaviour, and missing a single weight out of a hundred during extraction does not necessarily imply that the extracted model will behave correctly on 99% of the input space. While [8] introduced the notion of $(\epsilon,\delta)$-functional equivalence to capture such nuances, we propose a complementary metric—coverage—designed to quantify the proportion of the model that was not accurately recovered. To compute this metric, we randomly sample $10^{6}$ random input points and count how many activate the missed components in the target layer, or activate both a neuron in the previous layer and a neuron in the target layer connected by a missed weight. These input points, where the extracted model diverges from the target due to incomplete recovery, define the unrecovered space. The remaining inputs define the recovered space. The ratio of the recovered space to the total sampled input space gives the coverage score. When evaluating layer coverage, the coverage is measured with respect to a single layer. For model coverage, coverage is measured across the entire model.
Coverage serves two key purposes. First, it enables us to make a clear distinction between errors coming from imprecisions and errors coming from a lack of information, as it does not depend on the precision of the extraction. Second and foremost, coverage aims to capture the fact that for deep neural networks, we might not recover all the weights and yet still make an extraction correct on most of the input space. If we were to make an $(\epsilon,\delta)$ assessment of such an extraction, the results would be meaningless except for a $\delta$ corresponding to the coverage. Indeed, we would see a sudden increase in $\epsilon$ as soon as $\delta$ becomes greater than the coverage. $(\epsilon,\delta)$ assessments are still valuable on the recovered space. Our extraction achieves at least $(0.05,10^{-8})$ on the recovered part of each network (recall that unlike [8] we normalise $\epsilon$ in Definition 6). This means that if we recover each layer in this per-layer setting and recover the signs of the neurons, the output of the extracted network in the recovered space is within 5% of the target network’s output. We used $10^{8}$ random points in the recovered space to compute our $(\epsilon,\delta)$ values.
@@
Table 1: Results demonstrating the extraction of layers deeper than those recovered by state-of-the-art methods in [13]. Each layer is attacked independently using $3,000$ critical points. Layer 9 is linear and can therefore be trivially recovered using previously issued queries.
Model Signature Extraction L1 L2 L3 L4 L5 L6 L7 L8 L9 Entire Model
@@
Layer coverage $100\%$ $100\%$ $98.99\%$ $99.45\%$ $97.93\%$ $95.02\%$ $97.09\%$ $95.93\%$ $100\%$ $91.78\%$
CIFAR-10 Weights Recovered $100\%$ $100\%$ $90.63\%$ $98.44\%$ $93.75\%$ $92.19\%$ $95.31\%$ $93.75\%$ $100\%$ $99.91\%$
$-13072\text{-}8^{(8)}\text{-}$ 1 (I) Time $56$m48484848s $1$h33333333m $1$h42424242m $4$h43434343m $4$h25252525m $5$h2222m $6$h53535353m $2$h23232323m $0.01$s -
Queries $2^{23.55}$ $2^{26.15}$ $2^{26.15}$ $2^{27.10}$ $2^{27.04}$ $2^{27.17}$ $2^{27.56}$ $2^{26.45}$ 00 -

### v2

@@ 5.1 Evaluation Metrics
In practice, an attacker cannot query the entire input domain $\mathbb{R}^{d_{0}}$. Whether by design or due to constraints, extraction is restricted to a smaller region $S\subseteq\mathbb{R}^{d_{0}}$. On this space, which we require to be connected, some neurons do not have critical points and are thus always active (on) or always inactive (off). Always-off neurons can safely be ignored as they do not contribute to the network for inputs in $S$. Always-on neurons act as a fixed linear map on $S$. Under our goal of functional equivalence on $S$, this is naturally absorbed when recovering the final layer of $f$. There is no need to identify these neurons individually. If neuron $\eta^{(i)}_{k}$ is always off on $S$, then the $k$-th column of $A^{(i+1)}$ never contributes to the output on $S$, and recovering those weights is unnecessary. Many situations can cause a weight to never influence the network’s output on $S$. Thus, we introduce two definitions formalizing what the attacker is, in practice, aiming to recover.
Definition 7(effective architecture)
Let $f$ be a ReLU network with architecture $d_{0},d_{1},\dots,d_{r},d_{r+1}]$ and let $S\subseteq\mathbb{R}^{d_{0}}$. For $i=1,\dots,r$, let $n_{i}$ denote the number of neurons $\eta^{(i)}$ in layer $i$ whose pre-activation has a non-trivial zero set on $S$, i.e., $\{\,x\in S\mid\eta^{(i)}\!\circ F^{(i-1)}(x)=0\,\}\notin\{\emptyset,S\}$. Equivalently, $n_{i}$ counts the neurons that are neither always off nor always on over $S$. The effective architecture of $f$ on $S$ is then defined as $d_{0},\,n_{1},\,n_{2},\,\dots,\,n_{r},\,d_{r+1}]_{S}$.
Definition 8(effective weight)
Let $A^{(i)}=(a^{(i)}_{j,k})$ be the weight matrix from layer i1i-1 to layer $i$ and let $S\subseteq\mathbb{R}^{d_{0}}$. We say that the weight $a^{(i)}_{j,k}$ is effective on $S$ if it can influence the network’s output on $S$, that is, if there exists $x\in S$ such that
$\big(\eta^{(i-1)}_{k}\!\circ F^{(i-2)}(x)>0\big)\ \text{ and }\ \big(\eta^{(i)}_{j}\!\circ F^{(i-1)}(x)>0\big).$
We also say that $a^{(i)}_{j,k}$ is effective on $x$ as a shorthand for effective on $\{x\}\subseteq S$.
Now that we know what we aim to recover, i.e., effective weights in a network which behaves according to its effective architecture, we need metrics indicating the quality of the extraction. The number of effective weights recovered is not sufficient. Indeed, not all weights contribute equally to the network’s behaviour: failing to recover 1% of the weights does not generally mean that the extracted model will behave correctly on 99% of $S$. While [8] introduced the notion of $(\epsilon,\delta)$-functional equivalence to go beyond weights count, we propose a complementary metric, that we call coverage, to quantify the fraction of $S$ on which all effective weights are recovered.
00.10.20.30.40.50.60.70.80.911e-61e-41e-211 - $\delta$$\epsilon$($\log_{10}$)Model IModel IIModel Coverage
Figure 8: $(\epsilon,\delta)$ analysis of our end-to-end extractions of Models I and $I$. As the reduced-space proportion $\delta$ increases, the maximum output error $\epsilon$ decreases. The relationship is initially very sharp but smooths out after a transition phase, which aligns with the model coverage.
Definition 9(coverage)
Let $S^{\prime}\subseteq S\subseteq\mathbb{R}^{d_{0}}$ be the set of all $x$ for which no unrecovered weight of the model (resp. layer) is effective. The model (resp. layer) coverage is the proportion of points of $S$ that are also in $S^{\prime}$. We refer to $S^{\prime}$ as the covered space.
Representing volumes, coverage is estimated by sending a large set of random input points through the network and checking how many of them rely on unrecovered weights. The main purpose of coverage is to distinguish errors due to missing information (unrecovered but effective weights) from those due to numerical imprecision, since coverage is independent of the extraction’s numerical precision. Unlike an $(\epsilon,\delta)$ assessment (see Def. 6), which requires choosing a tolerance $\epsilon$ and whose meaning depends on the task/scale (the same $\epsilon$ can have a very different interpretation across networks), coverage is parameter-free and task-agnostic. Setting $\delta$ to 11- coverage and computing $\epsilon$ over the covered space provides a cleaner view of precision-improvement strategies: in practice, $\epsilon$ stabilizes quickly once $1-\delta$ falls below the coverage (Fig. 8).
5.2 Unrecovered Effective Weights
Our aim is to increase the coverage by recovering as many effective weights as possible given an attack cost. Let’s discuss the effective weights we do not recover. They fall in two distinct categories.
Definition 10(taxonomy of unrecovered weights)
For a given weight $a^{(i)}_{j,k}$, let $S_{(+,+)}\subseteq S\subseteq\mathbb{R}^{d_{0}}$ be the set of input points activating the two neurons connected by $a^{(i)}_{j,k}$:
$S_{(+,+)}=\{x\in S|(\eta_{k}^{(i-1)}\circ F^{(i-2)}(x)>0)\land(\eta_{j}^{(i)}\cdot F^{(i-1)}(x)>0)\}$
Similarly, let $S_{(+,-)}\subseteq S$ be the set of input points activating, out of the two neurons connected by $a^{(i)}_{j,k}$, only the neuron on layer i1i-1:
$S_{(+,-)}=\{x\in S|(\eta_{k}^{(i-1)}\circ F^{(i-2)}(x)>0)\land(\eta_{j}^{(i)}\cdot F^{(i-1)}(x)\leq 0)\}$
We define the $(+,+)$ (resp. (+,)(+,-)) activation of $a^{(i)}_{j,k}$ as the proportion of points of $S$ that are also in $S_{(+,+)}$ (resp. $S_{(+,-)}$).
For example, a weight has a $(+,+)$ activation equal to 0 if and only if $S_{(+,+)}=\emptyset$. Based on $(+,+)$ and (+,)(+,-) activations, we give a classification of unrecovered weights as follows:
Taxonomy of unrecovered weights $\mathbf{(+,+)}$ $\mathbf{(+,-)}$
Non-effective weights 0 $\geq 0$
Effective weights unreachable weights >0>0 0
query-intensive weights >0>0 >0>0
Query-intensive weights. To recover a weight, we must hit some specific hyperplanes. The budget for the attack is limited and thus we cannot grid the whole $S$ with search lines. The targeted search, one of the few steps in the attack which we do not improve, has its limitations too. Therefore it is expected that we miss some effective weights. We refer to those weights as query-intensive weights. The number of critical points of a neuron we gather is not clearly correlated to how close it is to being always on/off. Despite this lack of correlation at the neuron level, we observe that in general increasing the number of critical points used reduces significantly the number of query intensive weights, see Table 3.
Table 3: Taxonomy of unrecovered weights in all extracted models, under the assumption that all preceding layers have been perfectly extracted, for different attack costs. Model I having a much larger input size, we ran into memory limitations when performing the attack with 6,0006{,}000 critical points.
Models I II III
Critical points used 3,0003{,}000 3,0003{,}000 6,0006{,}000 3,0003{,}000 6,0006{,}000
non-effective weights 98 104 102 130 96
unreachable weights 5 8 8 0 0
query-intensive weights 18 11 5 25 7
Model Coverage $91.75\%$ $74.12\%$ $74.78\%$ $71.35\%$ $79.53\%$
Unreachable weights. Suppose that all critical points of a neuron $\eta_{j}^{(i)}$ lie on the inactive side of a neuron $\eta_{k}^{(i-1)}$. In this case, the contribution of $\eta_{k}^{(i-1)}$ to the output of $\eta_{j}^{(i)}$ is always zero during extraction, leaving us with no information about the corresponding weight $a_{j,k}^{(i)}$. If no region of $S$ exists where both neurons are active (as illustrated in Fig. 9(a)), then $a_{j,k}^{(i)}$ is never activated and can be safely treated as non-effective. However, if a region does exist where both neurons are active (depicted in grey in Fig. 9(b)), then $a_{j,k}^{(i)}$ is effective, contributes to the output of the network for inputs in this region, and should ideally be recovered. In this case, we say that $a_{j,k}^{(i)}$ is an unreachable weight. Increasing the number of critical points does not diminish the number of unreachable weights, see Table 3.
$\eta_{j}^{(i)}$$\eta_{k}^{(i-1)}$++
(a) $a_{j,k}^{(i)}$ is non-effective.
$\eta_{j}^{(i)}$$\eta_{k}^{(i-1)}$++
(b) $a_{j,k}^{(i)}$ is unreachable (and so effective).
Figure 9: The arrangement of hyperplanes can cause some effective weights to not be recoverable by the current extraction strategies. We call them unreachable.
6 Experimental Results
All our experiments were conducted on an Intel Core i7-14700F CPU using networks trained on either the MNIST or CIFAR-10 datasets, two widely used benchmarking datasets in computer vision. They have also been commonly used in related works [6, 8, 12] to evaluate model extraction methods. MNIST consists of 28 $\times$ 28 pixel grayscale images of handwritten digits, divided into ten classes (“0” through “9”) [16]. In contrast, CIFAR-10 contains 32 $\times$ 32 pixel RGB images of real-world objects across ten classes (e.g., airplane, automobile, bird, cat, deer, dog, frog, horse, ship, truck) [15]. The search space is the box centered at the origin of radius $10^{3}$ so as to keep the effective architecture close to the original.
6.1 Comparison with the State-of-the-Art
On a $784\!-\!16^{(8)}\!-\!1$ MNIST network, the state-of-the-art extraction of [12] fails to recover any hidden layer beyond the third, even under the optimal assumption that all previous layers have been perfectly extracted. On the same architecture, with the same assumptions, our attack recovers at least $93\%$ of the effective weights in each deeper layer. We further evaluate our method on other architectures and observe similar performance. The results are given in Table 4.
@@ 6.1 Comparison with the State-of-the-Art
Table 4: Results demonstrating the weight extraction of layers deeper than those recovered by state-of-the-art methods in [12]. Each layer is attacked independently using $3,000$ critical points. Layer 9 is linear and can therefore be trivially recovered using previously issued queries.
Model Signature L1 L2 L3 L4 L5 L6 L7 L8 L9 Entire
Extraction Model
@@
CIFAR-10 Effective Neurons 88 88 88 88 66 88 88 66 $10$ -
$3072\text{-}8^{(8)}\text{-}10$ Eff. Weights Recovered $\% 100100\%$ $90.32\%$ $98.36\%$ $9.74\% 88$$8.10\%$ $94.92\%$ $90.70\%$ $100\%$ $99.91\%$
(I) Time 5656m4848s 11h3333m 11h4242m 44h4343m 44h2525m 55h22m 66h5353m 22h2323m $0.01$s 2727h3838m
Queries $^{23.55} 22$$^{26.15}$ $^{26.15} 22$$^{27.10}$ $^{27.04} 22$$^{27.16}$ $^{27.56} 22$$^{26.45}$ 0 $2^{29.72}$
MNIST Effective Neurons 88 77 88 77 66 77 88 66 11 -
$784\text{-}8^{(8)}\text{-}1$ Eff. Weights Recovered $\% 100100\%$ $\% 100100\%$ $85.37\%$ $90.91\%$ $88.24\%$ $92.86\%$ $100\%$ $99.71\%$
(II) Time 33m2424s 77m55s 77m3737s 1010m4141s 1111m4646s 2828m2929s 2020m3838s 3636m4646s $0.01$s 22h66m
Queries $^{21.59} 22$$^{24.21}$ $^{24.22} 22$$^{24.24}$ $^{24.26} 22$$^{25.35}$ $^{24.96} 22$$^{25.61}$ 0 $2^{27.64}$
MNIST Effective Neurons $16$ $16$ $16$ $16$ $16$ $15$ $16$ $12$ 11 -
$784\text{-}16^{(8)}\text{-}1$ Eff. Weights Recovered $\% 100100\%$ $100\%$ $9.61\% 99$$9.21\%$ $97.01\%$ $98.64\%$ $93.78\%$ $100\%$ $99.82\%$
(III) Time 66m1717s 2626m1313s 2828m5757s 3939m3333s 11h3939m 22h1313m 11h4848m 33h5656m $0.01$s 1111h1717m
Queries $^{21.60} 22$$^{24.22}$ $^{24.24} 22$$^{24.66}$ $^{25.63} 22$$^{26.02}$ $^{25.78} 22$$^{26.93}$ 0 $2^{28.48}$
Foerster et al. [12]
MNIST All Weights $\% 100100\%$ - $\geq 37.50\%^{\dagger}$ - - - $\geq 0\%^{\dagger}$ $100\%$ -
$784\text{-}16^{(8)}\text{-}1$ Time 22h4646m 77m5050s - >36 h - - - >36 h $0.01$s -
Queries $2^{22.36}$ $2^{19.01}$ - - - - - - 0 -
L stands for layer. L1 is attacked with $500$ critical points. ’Eff.’ is short for effective. Foerster et al. recover 6/16 neurons on layer 4 and 0/16 neurons on layer 8, running the attack for 36 hours before stopping.
The taxonomy of these extraction results, as well as those obtained using 6,0006{,}000 critical points, is given above in Table 3. These results confirm our expectations. First, despite a large $S$, we observe a large amount of non-effective weights, about as much as one or two hidden layers. This shows that previously used metrics in the literature, notably the numbers of total weights recovered, can be misleading. Second, increasing the attack cost has the intended effects on the quality of the extraction: query-intensive weights diminish unlike unreachable weights. By doubling the number of critical points, Model II’s coverage barely changed from 74.12% to 74.78% while Model III’s, which does not have any unreachable active weights, changed more substantially from 71.35% to 79.53%. For both models only 0.9% additional weights are recovered.
6.2 Results of the End-to-End Attack
We give in Table 5 the results for the end-to-end extraction. For sign recovery, we employ the extension strategy. In model III, the extraction process terminates at layer 7, where only 12 of 16 neuron signatures are recovered with low precision, thereby preventing sign extraction and the subsequent signature extraction at layer 8. Note how, contrary to the previous extraction, the proportion of effective weights recovered plunges from layer five onwards.
Table 5: End-to-end attack results using $3,000$ critical points for all but the first layer. Layer 9 is linear and can therefore be trivially recovered using previously made queries.
Model End-to-end L1 L2 L3 L4 L5 L6 L7 L8 L9 Entire
Extraction Model
Effective Neurons 88 88 88 88 66 88 88 66 $10$ -
Layer coverage $\% 100100\%$ $98.97\%$ $75.11\%$ $98.43\%$ $95.06\%$ $7.12\% 99$$9.06\%$ $100\%$ $73.47\%$
CIFAR-10 Eff. Weights Recovered $\% 100100\%$ $90.32\%$ $91.67\%$ $92.86\%$ $90.48\%$ $94.92\%$ $93.02\%$ $100\%$ $99.90\%$
$3072\text{-}8^{(8)}\text{-}10$ Time for Signature 11h1414m 11h3535m 11h3838m 22h3838m 55h3535m 33h1919m 11h4141m 11h3838m $0.01$s 1919h1818m
(I) Queries for Signature $^{24.56} 22$$^{26.15}$ $^{26.15} 22$$^{26.58}$ $^{27.4} 22$$^{26.76}$ $^{26.18} 22$$^{26.22}$ 0 $2^{29.42}$
Time for Signs $1.69$s 2121s $43$s 22m11s 22m2525s 22m4343s 22m3030s 11m5151s $0.01$s 1212m3636s
Queries for Signs $^{24.17} 22$$^{22.85}$ $^{24.19} 22$$^{25.52}$ $2^{19.87}$ $^{19.64} 22$$^{26.83}$ $2^{19}$ 0 $2^{27.68}$
Effective Neurons 88 77 88 77 66 77 88 66 11 -
Layer coverage $\% 100100\%$ $\% 100100\%$ $\% 93.7593.75\%$ $76.21\%$ $88.50\%$ $100\%$ $74.82\%$
MNIST Eff. Weights Recovered $\% 100100\%$ $\% 100100\%$ $85.37\%$ $87.5\%$ $86.27\%$ $92.86\%$ $100\%$ $99.68\%$
$784\text{-}8^{(8)}\text{-}1$ Time for Signature 1212m88s 77m55s 77m4646s 1111m1313s 1313m2323s 1010m2121s 2222m77s 1919m4848s $0.01$s 11h4343m
(II) Queries for Signature $^{24.36} 22$$^{24.21}$ $^{24.22} 22$$^{24.24}$ $^{24.25} 22$$^{24.27}$ $^{24.97} 22$$^{24.78}$ 0 $2^{27.44}$
Time for Signs $0.19$s $0.85$s $47$s $1.46$s $1.33$s 11m3535s 11m2424s $12$s $0.01$s 44m22s
Queries for Signs $2^{20.23}$ $^{16.79} 22$$^{24.26}$ $2^{20.32}$ $^{15.68} 22$$^{25.24}$ $^{25.48} 22$$^{22.4}$ 0 $2^{26.77}$
Effective Neurons $16$ $16$ $16$ $16$ $16$ $15$ $16$ - - -
Layer coverage $\% 100100\%$ $100\%$ $98.78\%$ $88.28\%$ $87.69\%$ $28.09\%$ - - -
MNIST Eff. Weights Recovered $\% 100100\%$ $100\%$ $99.61\%$ $91.63\%$ $89.57\%$ $76.39\%$ - - -
$784\text{-}16^{(8)}\text{-}1$ Time for Signature 2121m3535s 2424m99s 2424m4747s 3939m1616s 11h5959m 11h1111m 22h4444m - - -
(III) Queries for Signature $^{23.69} 22$$^{24.31}$ $^{24.27} 22$$^{24.69}$ $^{26.01} 22$$^{25.1}$ $2^{25.63}$ - - -
Time for Signs $0.73$s 77m4444s 77m4949s 55m3232s 88m2222s 77m1111s 77m3838s - - -
Queries for Signs $^{20.23} 22$$^{24.7}$ $^{25.31} 22$$^{25.64}$ $^{25.68} 22$$^{26.4}$ $2^{26.43}$ - - -
L stands for layer. L1 is attacked with $500$ critical points.
The sudden drop in coverage in layer 4 of Model I is caused by a weight with a $(+,+)$ activation of $23\%$ of an almost-always-on neuron ($98\%$ active). Either making more queries, or slightly reducing the search space $S$ should increase substantially the coverage. The drop in coverage in layer 7 of Model II is caused by a particularly active unreachable weight with a $(+,+)$ activation of $11\%$. The $(\epsilon,\delta)$ analysis of the extraction of Model I and II is given in Fig. 8. We reach a $(3.64\times 10^{-4},0.26)$ extraction for Model I and a $(2.96\times 10^{-5},0.25)$ extraction for Model II. The choice of $\delta$ values is guided by their respective model coverage.
7 Conclusion
In this work, we carried out an in-depth analysis of the extraction approach introduced in [8], reused in several follow-up works [6, 7, 12], and identified critical limitations that prevented recovering weights beyond the third hidden layer. We then proposed improvements to the sign-recovery process by cleverly combining two methods introduced from Canales-Martínez et al. [6]. Finally, we introduced several techniques that together improve the numerical precision of the extracted signature. Taken as a whole, these advances permit, for the first time, polynomial end-to-end extraction of neural networks of significant depth and open the way to practical attacks on larger architectures.
Limitations. The main limitation we faced to going much deeper with the non end-to-end attack is the influence of deeper layers when targetting small-width networks, and the cost of the attack when targetting large-width networks. For the end-to-end attack, floating-point imprecisions remain the main limitation. Addressing these issues is a promising direction for future work.
Acknowledgements. The authors with NTU affiliation are supported by the Singapore NRF-NRFI08-2022-0013 grant. Christina Boura is partially supported by the France 2030 program under grant agreement No. ANR-22-PECY-0010 Cryptanalyse. Haolin Liu is also supported by the National Natural Science Foundation of China (No. 62372294) and the China Scholarship Council (No. 202406230017).
Appendix 0.A Example of a Small Neural Network
We illustrate here the definitions introduced in Section 2.1. Consider the 23212-3-2-1 neural network given by the following layers $A^{(i)}$, $b^{(i)}$: